Calculate amount of heat required to convert 1 kg of steam from hundred degree Celsius to 200 degrees celsius steam

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Specific heat of steam=1.996kJ/kg-K

Hence required heat=mC_s\Delta T=1kg\times 1.996kJ/kg-K\times (473-373)K=199.6kJ

Hope it helps. Regards.